\newpage
\begin{thebibliography}{99}
\bibitem{pp} P. Peld\'{a}n, "Large Diffeomorphisms in (2+1)-quantum Gravity
on the Torus", Penn State preprint 95/1-1, gr-qc/9501020
\bibitem{carlip1} S. Carlip, "(2+1)-dimensional Chern-Simons gravity as a Dirac
square
root", Phys. Rev. {\bf D 45} (1992) 3584 \\
S. Carlip, "Observables, Gauge Invariance and Time in (2+1)-Dimensional
Quantum Gravity", Phys. Rev. {\bf D42} (1990) 2647
\bibitem{ashtekar1} A. Ashtekar, V. Husain, C. Rovelli, J. Samuel and L.
Smolin, "2+1
quantum gravity as a toy model for the 3+1 theory", Class. Quant. Grav. {\bf 6}
(1989)
L185 \\
A. Ashtekar and R. Loll, "New Loop Representations for 2+1 Gravity",
Class. Quant. Grav. {\bf 11} (1994) 2417, gr-qc/9405031
\bibitem{louko} J. Louko and D. M. Marolf, "Solution Space of 2+1 Gravity on
$\car
\times T^2$ in Witten Connection Formulation", Class. Quant. Grav. {\bf
11} (1994) 311, gr-qc/9308018
\bibitem{woodhouse} N. Woodhouse, "Geometric Quantization", (Clarendon Press,
Oxford, 1980)\\
A. Ashtekar and invited contributors, "New Perspectives in Canonical Gravity",
 (Lecture Notes),(Bibliopolis, Napoli, Italy, 1988)
\bibitem{ADM} A. Hosoya and K. Nakao, (2+1)-Dimensional Pure Gravity for an
Arbitrary
Closed Initial Surface", Class. Quant. Grav. {\bf 7} (1990) 163 \\
Y. Fujiwara and J. Soda, "Teichm\"{u}ller Motion of (2+1)-Dimensional Gravity
with the
Cosmological Constant", Progr. Theor. Phys. {\bf 83} (1990) 733 \\
V. Moncrief, "How Solvable is (2+1)-Dimensional Einstein Gravity", J. Math.
Phys. {\bf
31} (1990) 2978, \\ "Reduction of the Einstein Equations in 2+1 Dimensions to a
Hamiltonian System over Teichm\"{u}ller Space", J. Math. Phys. {\bf 30} (1989)
2907
 A. Hosoya and K. Nakao, "(2+1)-Dimensional Quantum Gravity", Prog.
Theor. Phys. {\bf 84} (1990) 739
\bibitem{carlip2} S. Carlip, "Six Ways to Quantize (2+1)-dimensional Gravity",
gr-qc/9305020
\bibitem{tate} A. Ashtekar and R. S. Tate, "An Algebraic Extension of Dirac
Quantization: Examples", J. Math. Phys. {\bf 35} (1994) 6434, gr-qc/9405073
\bibitem{riemm} T. Miyake, "Modular Forms", (Springer-Verlag,
Berlin Heidelberg, 1989)\\
 H. Maass, Lecture Notes: "Lectures on Modular Functions of One Complex
Variable", (Tata Institute of Fundamental Research, Bombay, 1964) \\
A. Terras, "Harmonic Analysis on Symmetric Spaces and Applications I",
(Springer Verlag, New York, 1985)
\bibitem{marolf} D. M. Marolf, "Loop Representations for 2+1 Gravity on a
Torus",
Class. Quant. Grav. {\bf 10} (1993) 2625, gr-qc/9303019

\end{thebibliography}

\documentstyle[12pt]{article}
\newcommand{\bea}{\begin{eqnarray}}
\newcommand{\be}{\begin{equation}}
\newcommand{\eea}{\end{eqnarray}}
\newcommand{\ee}{\end{equation}}

\def\l{\lambda}
\def\m{\mu}

\def\G{\Gamma}
\def\J{\Psi}

\def\cc{{\cal C}}

\def\cm{{\cal M}}

\def\cq{{\cal Q}}
\def\car{{\cal R}}

\def\cz{{\cal Z}}

\begin{document}
\pagestyle{plain} \large \noindent . \vspace*{5mm} \\
CGPG-95/4-2\\gr-qc/9504014 \\April 1995
\begin{center} \LARGE \bf \vspace*{5mm}
A Modular Invariant Quantum Theory From the Connection Formulation of
(2+1)-Gravity on the Torus\\
\vspace*{10mm} \large \bf
 Peter Peld\'{a}n
\footnote{Email address: peldan@phys.psu.edu}\\
\vspace*{5mm} \large
 Center for Gravitational Physics and Geometry\\
Physics Department\\
The Pennsylvania State University, University Park, PA 16802, USA\\
\end{center}
\begin{abstract}
By choosing an unconventional polarization of the connection phase space in
(2+1)-gravity on the torus, a modular invariant quantum theory is constructed.
Unitary equivalence to the ADM-quantization is shown.
\\ PACS numbers: 04.60.-m, 04.60.Ds, 04.60.Kz \end{abstract}
\normalsize
It has recently been shown \cite{pp} that there are severe obstructions to
constructing a quantum theory for (2+1)-dimensional gravity on a manifold $\car
\times T^2$ in the connection representation if we treat all diffeomorphisms
as gauge. That is, we require physical states in our Hilbert space to be
invariant -- or at most transform according to a one-dimensional unitary
representation of the diffeomorphism group -- under all diffeomorphisms. In
\cite{pp} only the space-like sector \cite{carlip1, ashtekar1, louko} was
considered, but
the result can easily be transferred to the time-like sector as well.

The purpose of this Letter is to show that this problem can be solved
straightforwardly by choosing a different polarization of the phase space
\cite{woodhouse}.
The polarization chosen here leads automatically to the Heisenberg picture of
the ADM-formulation \cite{ADM, carlip2}.

The reason why we get such severe problems in trying to construct a quantum
theory in the connection representation with all diffeomorphisms treated as
gauge, is that the group of large diffeomorphisms has a very bad behavior on
the reduced configuration space; the group of large diffeomorphisms on the
torus does not act properly discontinuous in the space of gauge equivalent
classes of flat $SO(1,2)$ connections on the torus. However, we know that the
action of this group is nicely behaved in the phase space -- except for a
region of measure zero -- and we should therefore be able to quantize this
formulation with the use of a different polarization.\\ \\
In the space-like sector of (2+1)-gravity on the torus, we have, after
solving all constraints and gauge fixing all local symmetries, the following
theory \cite{carlip1, ashtekar1, louko}: the reduced phase space is the
cotangent
bundle over the reduced configuration space $\cq =(\car ^2-\{0,0\})/\cz _2$
coordinatized by the
coordinates $\m $ and $\l $, where $\cz _2$ acts in
$\car ^2$ as $\{\m , \l \}\rightarrow \{-\m ,-\l \}$.
The canonically conjugate momenta are denoted $a$ and $b$ and we have the two
fundamental Poisson brackets: $\{ \m ,a\}=\frac{1}{2}$ and $\{ \l
,b\}=-\frac{1}{2}$. Note that the $\cz _2$ identification in the phase space
reads; $(\m ,\l ,a,b)\sim -(\m ,\l ,a,b)$.
 The Hamiltonian is identically zero. Since all local
symmetries have been taken care of, the only remaining
gauge redundancies are the large diffeomorphisms and the large $O(1,2)$
transformations. The large diffeomorhisms on the torus form a
group -- the modular group, $PSL(2,\cz )=SL(2,\cz)/\{{\bf 1},{\bf -1}\}$ --
which is generated by
\bea S: && (a,\l )\rightarrow (b,\m ),\hspace{10mm} (b,\m )\rightarrow
(-a,-\l ) \label{S}\\
T:&& (a,\l )\rightarrow (a,\l ), \hspace{10mm} (b,\m ) \rightarrow (b+a,\m +
\l ) \label{T}\eea

while the remaining large $O(1,2)$ part is isomorphic to $\cz _2$ and
generated by
\be O:\hspace{5mm}(a,\l )\rightarrow (a,-\l ),\hspace{10mm} (b,\m
)\rightarrow (b,-\m ) \label{otransf} \ee
To quantize this system, I will follow the algebraic quantization program
outlined in \cite{tate} while treating the above large symmetries as gauge.
Alternatively, one may see the quantization below as a straightforward
application of standard geometrical quantization \cite{woodhouse} with a
special polarization of the phase space.
Instead of quantizing the fundamental Poisson-algebra in the vector
space of complex valued functions in $\cq $, I choose to quantize the
Poisson-algebra of the phase space functions\footnote{The reasons for this
choice is partly that these functions are good coordinates on phase space --
the Jacobian is everywhere non zero -- and partly that the modular group now
has a nice action in $m,\bar{m}$ space except for the region $m-\bar{m}=0$
$\Leftrightarrow $ $a\m - b\l =0$.}
\bea m:=\frac{(b+ i\m )}{(a+ i \l )}&& \bar{m}:=\frac{(b-i \m )}{(a- i \l )}
\label{m}\\
p:=-i (a- i \l )^2 && \bar{p}:=i (a+ i \l )^2 \label{p}\eea
in the vector space of complex valued functions of $m$ and $\bar{m}$. Note
that these functions automatically takes care of the redundancy
identification $(\m , \l ,a,b)\sim -(\m , \l ,a,b)$, while the large
diffeomorphisms and $O(1,2)$ transformations become
\bea S:&& m\rightarrow -\frac{1}{m},\hspace{10mm} p\rightarrow m^2p
\label{S2}\\
T: && m\rightarrow m+1,\hspace{10mm} p\rightarrow p \label{T2} \\
O: && m\rightarrow \bar{m},\hspace{10mm} p\rightarrow -\bar{p}\label{O2} \eea
and similarly for their complex conjugate functions.
Explicitly, the classical Poisson-algebra is
\bea && \{m,\bar{m}\}=\{m,p\}=\{\bar{m},\bar{p}\}=\{p,\bar{p}\}=0 \\
&& \{m,\bar{p}\}=\{\bar{m},p\}=-2 \eea
which is easily quantized by the representation
\be \hat{m}=m,\hspace{5mm}\hat{\bar{m}}=\bar{m},\hspace{5mm}
\hat{p}=2i\frac{\partial}{\partial \bar{m}} + \frac{1}{m_2}, \hspace{5mm}
\hat{\bar{p}}:=2i
\frac{\partial}{\partial m} - \frac{1}{m_2} \ee
where $m=m_1 + im_2$. (The representation of $\hat{p}$ and $\hat{\bar{p}}$ is
chosen such that formally $\hat{p}^\dagger = \hat{\bar{p}}$ w.r.t the
inner-product
defined below.)

Next we need to find the physical states, {\it i.e.} complex valued functions
of $m$ and $\bar{m}$ that are invariant under the remaining "gauge
transformations" (\ref{S2}), (\ref{T2}) and (\ref{O2}). Since we know that
the quotient space ${\bf \cc} /PSL(2,\cz )/\cz _2 $ simply is the torus moduli
space\footnote{Actually, to be completely honest, this is only true
for the complex plane minus the real axis. However, since the real axis is a
region of measure zero in the complex plane, I will simply neglect it.},
$\cm $
 \cite{riemm}, we automatically know that our physical states are given by
functions on this space. Then, we need an inner-product, and since there
exist a unique -- up to scaling -- modular invariant measure \cite{riemm} on
the moduli space, the inner product is naturally chosen to be

\be <\J _1| \J _2 >:=\int _{\cm } \frac{dm_1 dm_2}{m_2^2}\bar{\J }_1\J _2
 \ee
Explicitly, the moduli space is $\cm $: $|m|\geq 1$, $m_2 >0$ and
$|m_1|\leq \frac{1}{2}$.

To complete the quantization one should now find
a complete set of $PSL(2,\cz )\otimes \cz _2$ invariant physical observables,
check that they are
self adjoint in the Hilbert space, and check if their algebra is irreducibly
represented in the Hilbert space. I will not address these issues here.

Thus, this completes the construction of a non-trivial Hilbert space carrying
a one-dimensional unitary representation of the modular group.\\ \\
Now, how is this quantum theory related to the ADM-quantization? Just by
comparing the definitions (\ref{m})-(\ref{p}) to the time dependent canonical
transformation between half the phase space\footnote{Explicitly we have $\G
_{ADM} \cong \G _{conn}/\cz _2$, where $\G $ denote the phase space and $\cz
_2$ acts in $\G _{conn}$ as in eq. (\ref{otransf}).} of the connection
formulation and
the ADM-formulation \cite{carlip1}, one immediately sees that the functions
$m$, $\bar{m}$, $p$ and $\bar{p}$ are nothing more than the ADM-variables at
a specific instant of ADM-time. Thus, the above
quantum theory is simply the Heisenberg picture of the ADM
Schr\"{o}dinger picture, and unitary equivalence follows directly.\\ \\
Finally we may note that although we have managed to construct a non trivial
modular invariant theory starting from the connection formulation, our
quantum theory is not a connection representation \cite{ashtekar1, marolf}.
This immediately implies that we do not have any obvious way to relate this
quantum theory to the so called loop-representation \cite{ashtekar1, marolf}.
 Furthermore,
it seems very unlikely that the above construction should have any chance of
working in the time-like sector where the modular group does not act properly
discontinuous even in the full phase space. Thus the construction of a non
trivial
modular invariant connection representation still remains as an open
problem.\\ \\
{\bf Acknowledgements}\\
I thank Abhay Ashtekar for suggesting solving the problem of ref.\cite{pp} by
choosing a different polarization. I also thank Don Marolf for discussions.

\end{document}